### v3

# Live-SWE-agent: Can Software Engineering Agents Self-Evolve on the Fly?

Abstract
Large Language Models (LLMs) are reshaping almost all industries, including software engineering. In recent years, a number of LLM agents have been proposed to solve real-world software problems. Such software agents are typically equipped with a suite of coding tools and can autonomously decide the next actions to form complete trajectories to solve end-to-end software tasks. While promising, they typically require dedicated design and may still be suboptimal, since it can be extremely challenging and costly to exhaust the entire agent scaffold design space. Recognizing that software agents are inherently software themselves that can be further refined/modified, researchers have proposed a number of self-improving software agents recently, including the Darwin-Gödel Machine (DGM). Meanwhile, such self-improving agents require costly offline training on specific benchmarks and may not generalize well across different LLMs or benchmarks. In this paper, we propose Live-SWE-agent, the first live software agent that can autonomously and continuously evolve itself on-the-fly during runtime when solving real-world software problems. More specifically, Live-SWE-agent starts with the most basic agent scaffold with only access to bash tools (e.g., mini-SWE-agent), and autonomously evolves its own scaffold implementation while solving real-world software problems. Our evaluation on the widely studied SWE-bench Verified benchmark shows that Live-SWE-agent can achieve an impressive solve rate of 77.4% without test-time scaling, outperforming all existing software agents, including the best proprietary solution. Moreover, Live-SWE-agent outperforms state-of-the-art manually crafted software agents on the recent SWE-Bench Pro benchmark, achieving the best-known solve rate of 45.8%. More details at: [Uncaptioned image] https://github.com/OpenAutoCoder/live-swe-agent
00footnotetext: Work done as a research intern at the University of Illinois Urbana-Champaign.
*Customized SWE-agent  (kimik2thinking)   **JSON-based apply_patch tool plus dynamic thinking time (gpt51)
Figure 1: SWE-bench Verified and SWE-Bench Pro results (single attempt w/o test-time scaling)
1 Introduction
Large Language Models (LLMs) have rapidly progressed from simple code autocompletion austin2021program; chen2021codex; li2022competition; xia2022less to interactive agents capable of navigating repositories, running tests, and submitting patches end‑to‑end yang2024sweagent; wang2024openhands; zhang2024autocoderover; xia2024agentless; carbonneaux2025cwm; liu2024large. Early conversational repair systems leveraged environment feedback to iteratively refine candidate fixes xia2024automated, while subsequent agentic frameworks (such as SWE-agent yang2024sweagent and OpenHands wang2024openhands) augmented LLMs with terminals, editors, and search, enabling multi‑step tool use on complex repositories. At the same time, complementary “more agentless” pipelines (e.g., Moatless moatless and Agentless xia2024agentless) argue that much of the perceived complexity in agent scaffold design can be replaced by dedicated prompting and workflow design.
Despite this progress, most existing agents have fixed designs and are limited in their static action space, where the scaffold implementation (including tools) is preset even when a task could benefit from more customization. In addition, manually designing an optimal software agent scaffold is extremely challenging and costly due to the infinite design space. As a result, more recently, the community has begun to explore self-evolving software agents zhang2025darwin; robeyns2025sica; wang2025huxley, which iteratively modify their own scaffold implementations and empirically validate each change using offline evaluation signals from coding benchmarks. However, these approaches add significant additional cost. For example, a single run of DGM on SWE-bench costs around \$22,000 according to the original paper zhang2025darwin. Furthermore, they rely heavily on offline evolution, where improvements are typically learned on certain benchmarks and then baked into a static agent. In this way, the learned agents may become specialized to the given benchmarks and underlying LLMs, and may not generalize well beyond that.
To bridge this gap, we introduce Live-SWE-agent, the first live, runtime self-evolving software engineering agent that expands and revises its own capabilities on the fly while working on a real-world issue. Our key insight is that software agents are themselves software systems, and modern LLM-based software agents already possess the intrinsic capability to extend or modify their own implementation at runtime. While our idea of enabling on-the-fly self-evolution applies to all parts of the agent scaffold implementation, as a first step, we focus primarily on tool creation, as it is one of the most essential parts of a software agent. Live-SWE-agent starts from a simplistic agent with only bash tool access (e.g., mini-SWE-agent yang2024sweagent). During the regular issue-solving loop, the agent can synthesize, modify, and execute custom tools such as editors, code search utilities, and domain-specific analyzers. A lightweight step‑reflection prompt repeatedly asks the agent whether creating or revising a tool would accelerate progress, thereby turning tooling into a first‑class decision alongside ordinary actions like running tests. This mechanism leaves the underlying scaffold unchanged, requires no offline training, and is agnostic to the underlying LLM. By elevating tool creation to an explicit, iterative decision point, we unlock this hidden on-the-fly self-improvement ability.
Live-SWE-agent addresses the limitations of existing research. First, with tool creation, the agent action space adapts to the problem at hand, producing task-relevant tools that precisely capture the needs to accomplish the current task. Furthermore, by shifting improvement from offline training to online evolution, Live-SWE-agent alleviates tedious scaffold engineering because new abilities emerge from the encountered issues themselves. Importantly, tool synthesis is not a one-shot preprocessing action but a continuous iterative process interleaved with problem solving. The agent can refine tools as their understanding of the failure mode evolves. Despite its minimal and simplistic design, Live-SWE-agent is generalizable to different agent scaffolds and LLMs, with state-of-the-art open results on software issue solving.
We have evaluated Live-SWE-agent on the widely used SWE-bench Verified benchmark sbv and the more challenging SWE-Bench Pro deng2025swe. Without any test‑time scaling, Live-SWE-agent achieves 77.4% resolve rate on Verified and 45.8% on Pro, surpassing state-of-the-art open‑source baselines and even the best commercial agent systems. Our ablations and tooling analyses reveal that (1) custom tool creation materially improves effectiveness (higher solve rates) with minimal overhead, (2) benefits persist across different state-of-the-art LLM backends, with better results on stronger models, demonstrating its promising future as LLMs’ capabilities rapidly evolve, and (3) synthesized tools include both generic utilities and task-aligned specializations that benefit issue-specific problem-solving.
In summary, we make the following contributions:
  • The first live software agent. We present Live-SWE-agent, the first live software agent that can autonomously self-evolve its own scaffold implementation on the fly when solving real-world issues, without any offline training or extra pipelines.
  • Minimal and general implementation. Our current implementation adopts a minimal and general design, where the agent starts from a minimal bash-only scaffold (e.g., mini-SWE-agent) and self-improves on the fly by creating general or customized tools. The design is compatible with any existing software agent loop or LLM with negligible overhead, has been publicly available at: https://github.com/OpenAutoCoder/live-swe-agent.
  • State-of-the-art performance. On SWE-bench Verified and SWE-Bench Pro, Live-SWE-agent reaches 77.4% and 45.8% solve rates (without any test-time scaling), respectively, outperforming all existing open-source agents and commercial systems at the time of writing. To our knowledge, our SWE-Bench Pro result is also the best reported to date.
  • Comprehensive analysis. We include detailed studies of when and why on-the-fly tool creation helps, how it improves efficiency, and how it compares to other designs like fixed-tool agents, workflow-based systems, and offline self-improving methods. Notably, compared to existing self-improving agents, we achieve much better performance (65.0% solve rate on the SWE-bench Verified-60 subset vs. DGM’s 53.3%) while being significantly less costly (no offline cost).
  • Unified leaderboard. For software tasks, recent LLMs are often benchmarked using manually engineered, proprietary agent scaffolds, making fair model comparison hard. Live-SWE-agent offers an open, unified, and powerful scaffold that enables genuinely fair, apples-to-apples comparison for future model releases. We are maintaining a leaderboard of recent models evaluated on real-world software tasks using Live-SWE-agent at: http://live-swe-agent.github.io
2 Approach
Figure 2: Overview of Live-SWE-agent
Live-SWE-agent is a live, self-evolving agent that improves and expands its capabilities on the fly while solving an issue. Our key insight is recognizing that agents themselves can be iteratively improved, just like the software issues they are designed to solve. The space in which an agent can evolve includes not only the tools it uses but also the underlying agent scaffold itself. In this work, we start from a simple agent scaffold and focus primarily on self-evolution for tool creation and usage, as tools represent one of the most critical components of an agent. We now describe how Live-SWE-agent provides a framework for LLMs to develop and use their own tools on the fly.
Figure 2 presents an overview of Live-SWE-agent. First, 1 we take in both the project codebase and the description of the issue to be solved. We then provide this information to the agent and initialize it with a set of tools. In the beginning, the agent may only have access to a limited set of tools (e.g., bash commands) and the goal of Live-SWE-agent is to allow the agent to generate and use its own tools on the fly while solving the issue. As the agent is solving the issue, at each step,it can choose to either 2 output a command (e.g., to use a tool) or 3 create a custom tool that can help it solve the issue. In Live-SWE-agent, we define a custom tool as a script that can be executed in the environment. This provides an intuitive and straightforward tool-usage interface for the agent, allowing it to output a command to use any custom tool. Next, different from existing approaches that 4 directly provide the environmental feedback output to the agent, 5 we specifically ask the agent to reflect upon the past steps and decide whether a tool should be created in the feedback message. This loop is repeated until the agent has submitted a solution 6 to the initial problem. Unlike prior agentic setups where the set of tools and possible actions available to the agent is fixed, Live-SWE-agent allows the agent to perform live self-evolution by creating and using custom tools on the fly.
2.1 On-the-fly Self Evolution
The key idea of Live-SWE-agent is to enable the agent self-improve by modifying its own scaffold, such as creating custom tools based on the problem and previous trajectories. To support tool creation, we apply several simple modifications to the initial prompt of the agent. Specifically, we provide instructions and examples illustrating how a tool should be created and used. More importantly, we indicate to the agent that: (1) the goal of any tools it creates should be to help it better solve the tasks and (2) the created tools can be for any purpose and do not need to be general.
In addition to introducing the ability for agents to create tools through the initial prompt, we also explicitly ask the agent to reflect on its past trajectory to determine if it should create any tools after each step. This is done by appending a simple reflection message after each environmental feedback. In our experiments, we found that this reflection process is necessary to remind the agents to design tools that are useful and specific for the particular issue.
It is important to point out that the modifications made to a regular agentic framework by Live-SWE-agent are extremely simple (See Appendix D for both the initial prompt and reflection message). Live-SWE-agent does not make any changes to the agentic loop, impose a particular workflow, or require any costly offline training. Instead, the focus is on allowing and extending the ability for agents to create their own custom tools at runtime to improve performance and reduce manual tool-creation effort. This allows Live-SWE-agent to be extremely general and applicable across a wide-range of different tasks, LLMs, and agent scaffolds. Furthermore, we also note that software agents, in essence, are also software. As such, they can be modified and updated on the fly by software agents (themselves) no different than any other software repository. In Live-SWE-agent, we leverage this insight for on-the-fly self-improvement by enabling the agents to create their own custom tools depending on the problem at hand. We next describe the custom tools agents create in more detail.
2.2 Custom Tool Synthesis
In Live-SWE-agent, we define a custom tool as a script that can be executed in the environment. This allows the agents to both easily create tools (by creating a script file) and use the created tools (by running the script with arguments). We believe this is a general and intuitive interface that is suitable for a wide range of tasks.
Example editing tool. Figure 3 shows an example of a custom tool created by the agent. This is an editing tool that allows the agent to edit a file by replacing, inserting, or deleting code. We see that it contains the necessary tool logic as well as clear instructions on how the tool can be used. Compared with bash commands which can potentially overwhelm the agent with many different arguments and flags, the tools that agents create themselves have a straightforward purpose and are easy to use. Furthermore, in the example, we see that the editing tool also provides relevant feedback messages such as indicating whether the replacement edit was successful or not. This feedback can be critically important to inform the next actions taken by the agents. On the other hand, a bash editing command like sed does not indicate the result of the edit such that a replacement operation where the string to be replaced does not exist will produce no warning message but still return a success code. As such, the agent may be misguided into thinking that the edit was successful when, in fact, no changes have been made to the file.
Live-SWE-agent also encourages agents to create tools that can improve the efficiency of their workflows. For example, a common step in solving an issue is to identify the locations of relevant code snippets in order to understand the root cause. In this case, a custom search tool can be useful for searching code within certain directories and displaying their surrounding context. While it is possible to achieve the same outcome of a custom search tool using a combination of different bash commands (e.g., grep, find, cat), the agent often needs to take multiple steps to perform the actual search, leading to increased context length and time required to solve an issue. By creating custom tools that can handle complex and multi-step tasks, Live-SWE-agent can improve both the effectiveness and efficiency of agents.
(a) Edit tool
(b) MARC file analyzer tool
Figure 3: Example custom tools
Example issue-specific tool. In addition to general tools, another benefit of Live-SWE-agent is its ability to create issue-specific tools. Figure 3 shows an example of a custom tool that is specifically tailored to a particular issue. In this example, the tool analyzes MARC files, a file format for publication or text records, and prints them in a human-readable format. The agent can create this tool and use it to display the content of relevant MARC files (including binary files) that serve as test cases, helping it better understand the issue and evaluate potential patches. Such functionality cannot be achieved easily using simple bash commands or even general-purpose tools. By enabling the agents to create arbitrary custom tools on the fly, Live-SWE-agent can generate specialized tools for individual problems, allowing it to solve issues more effectively.
A fair question to ask is why not generate these tools directly in the beginning? The reason is that tool creation, similar to manual problem solving, is also an iterative process. We need to understand the problem at hand and during the solving process identify the issues in order to come up with helpful tools in different scenarios. For example, in the MARC file example, it is not immediately obvious that we need to create a custom tool specifically to analyze MARC files. By creating all the tools in the beginning and applying this fixed set of tools during the entire process, we will lose the opportunity to design custom tools that are helpful in unique situations. Additionally, having access to all possible tools are not necessarily better as they can overwhelm and mislead the agent. Furthermore, the agent may often iterate on or modify the tools it has created which requires runtime modification abilities not available under a fixed tool setup. Live-SWE-agent provides the ability for agents to automatically synthesize customs tools on the fly without any additional overhead from heavy scaffold modifications or costly offline training updates.
3 Experimental Setup
Implementation. While Live-SWE-agent is general across different software agent scaffolds, we implement Live-SWE-agent on top of the popular mini-SWE-agent (yang2024sweagent) framework. As such, we retain the hyperparameters used in mini-SWE-agent by default (i.e., maximum step limit of 250 and maximum cost of \$3 per issue). We choose mini-SWE-agent to build on since it is not only simplistic (with $\sim$100 lines of code and only accessing bash commands) but also widely used. Unless otherwise stated, we use Claude 4.5 Sonnet (claude-sonnet-4-5-20250929(sonnet4.5) in our experiments and sample one patch per issue.
Datasets. We evaluate Live-SWE-agent on the popular SWE-bench Verified benchmark (sbv) containing 500 software development problems where the goal is to successfully modify the repository given a problem description. SWE-bench Verified is validated by human developers to ensure each problem description has sufficient amount of information to solve the issue. Additionally, we also evaluate on the recent SWE-Bench Pro (deng2025swe) benchmark, containing 731 publicly available problems, aimed to capture realistic, complex, and enterprise-level problems. Compared with SWE-bench Verified, SWE-Bench Pro contains more difficult problems across multiple repositories and programming languages.
Baselines. We compare Live-SWE-agent against representative state-of-the-art agentic approaches. For SWE-bench Verified, we compare against mini-SWE-agent as it is one of the most widely-used open-source agentic solutions on SWE-bench tasks with top leaderboard performance. Additionally, it is also a straightforward comparison as we directly build on top of the mini-SWE-agent framework. Furthermore, we compare with prior self-evolving agents: Self-Improving Coding Agent (SICA) (robeyns2025sica), Darwin-Gödel Machine (DGM) (zhang2025darwin), and Huxley-Gödel Machine (HGM) (wang2025huxley) on a subset of SWE-bench Verified problems. This subset of 60 SWE-bench Verified problems has been used by prior work (wang2025huxley) to specifically evaluate all three self-improving agent baselines. For SWE-Bench Pro, we compare against SWE-agent (yang2024sweagent) which is the top-performing approach on the SWE-Bench Pro leaderboard. For each baseline, we directly reuse their experimental results and report their performance, cost, as well as the backend LLM used whenever possible.
4 Evaluation
Table 1: Result on SWE-bench Verified
Tool LLM % Resolved Avg. \$ Cost
[Uncaptioned image] GPT-5-Mini 59.8% \$0.04
[Uncaptioned image] GPT-5 65.0% \$0.28
Claude 4.5 Sonnet
70.6% \$0.56
mini-SWE-agent (yang2024sweagent)
Gemini 3 Pro
74.2% \$0.46
[Uncaptioned image] GPT-5-Mini 63.0% \$0.05
[Uncaptioned image] GPT-5 68.4% \$0.27
Claude 4.5 Sonnet
75.4% \$0.68
Live-SWE-agent
Gemini 3 Pro
77.4% \$0.48
Table 2: Result on SWE-bench Verified-60
Tool LLM % Resolved Offline cost (hours)
Offline self- improving agents
SICA (robeyns2025sica) [Uncaptioned image] GPT-5-Mini 50.0% infinite loop
DGM (zhang2025darwin) [Uncaptioned image] GPT-5-Mini 53.3% 1231
HGM (wang2025huxley) [Uncaptioned image] GPT-5-Mini 56.7% 512
Live-SWE-agent [Uncaptioned image] GPT-5-Mini 65.0% 0
4.1 Main Results
SWE-bench Verified. Table 1 shows the results of Live-SWE-agent and prior agent approaches on SWE-bench Verified. We first observe that compared with mini-SWE-agent, across four different LLM backends, Live-SWE-agent consistently achieves a higher resolve rate. This demonstrates the improvement in performance enabled by allowing agents to create and use their own custom tools on the fly. Furthermore, Live-SWE-agent is able to achieve this with only a minimal increase in cost compared with the base mini-SWE-agent. In certain cases (e.g., in GPT-5), we even observed slight cost savings where the agent can improve the solving efficiency by replacing complex, multi-turn commands with customs tools that achieve the same functionality.
We also demonstrate the comparison between Live-SWE-agent and the best-performing agentic tools, including both state-of-the-art open-source solutions and proprietary commercial scaffolds, on SWE-bench Verified (Figure 1). We only report single-attempt results in Figure 1 without any test-time scaling to ensure a fair comparison. We observe that Live-SWE-agent using Gemini 3 Pro achieves a solve rate of 77.4% without test-time scaling, outperforming all existing agents on the SWE-bench Verified leaderboard [^1] https://www.swebench.com including state-of-the-art commercial solutions at the time of writing.
Furthermore, we perform a detailed comparison with three prior self-evolving agents on a subset of 60 SWE-bench Verified problems chosen by prior evaluation (wang2025huxley). Table 2 shows the resolve rate and the offline cost for SICA, DGM, HGM, and Live-SWE-agent on the subset of problems. We observe that Live-SWE-agent achieves the best performance, improving the resolve rate by 8.3 percentage points compared to previous best approach. We also see that previous self-evolving agents all require a heavy amount of offline training to evolve the base agent, costing more than 500 hours. Furthermore, prior self-evolving techniques produce a static agent used for all problems. On the other hand, Live-SWE-agent creates customs tools for each individual task, allowing it to adapt on the fly based on the problem and specific LLM used. Unlike the costly offline updates required by prior self-evolving agents, Live-SWE-agent instead adopts an online evolving approach to prompt the agent to generate custom tools on the fly to improve performance with minimal overhead.
Table 3: Result on SWE-Bench Pro
Tool LLM % Resolved Avg. \$ Cost
SWE-agent (yang2024sweagent)
Claude 4.5 Sonnet
43.6% -
Live-SWE-agent
Claude 4.5 Sonnet
45.8% \$0.73
SWE-Bench Pro. We also evaluate Live-SWE-agent on the public set of SWE-Bench Pro, containing 731 problems across 11 repositories and four programming languages (Python, Go, TypeScript, and JavaScript). Table 3 shows the performance of Live-SWE-agent compared to the state-of-the-art baseline of SWE-agent on SWE-Bench Pro deng2025swe. Different from mini-SWE-agent, which only has access to basic bash commands, SWE-agent provides handcrafted file viewing and editing tools and is the best performing approach on the SWE-Bench Pro. We choose Claude 4.5 Sonnet for SWE-agent because it ranks at the top of SWE-Bench Pro leaderboard [^2] https://scale.com/leaderboard/swe_bench_pro_public, as such, we also use it as the base LLM for Live-SWE-agent to enable a fair comparison. We observe that Live-SWE-agent is able to achieve better performance compared with SWE-agent, a specially designed agent implemented with close to 7,000 lines of code. We also compare with all other top-performing agents and LLMs in Figure 1 and observe that Live-SWE-agent is able to achieve the new state-of-the-art performance of 45.8% resolve rate on SWE-Bench Pro. This further demonstrates the superiority of the live scaffold design of Live-SWE-agent compared to existing agents that interact with a fixed set of manually-crafted tools.
4.2 Tools Analysis
(a) Tools in SWE-bench Verified
(b) Tools in SWE-Bench Pro
(c) Tools in SWE-Bench Pro
Figure 4: 2 dimensional t-SNE visualization of tools generated by Claude 4.5 Sonnet on SWE-bench Verified and SWE-Bench Pro. We label and display the embedding based on tool type (Figure 4(a)), repository name (Figure 4(b)), and programming language used in the repository (Figure 4(c)). Note that for Figure 4(b), we only label three repositories in the legend due to space considerations. The three repositories were chosen as they have representative distinct clusters.
Categories and variations of custom tools. We examine the custom tools created by Live-SWE-agent. Figure 4 shows the embedding visualizations using t-SNE (hinton2002stochastic) of the tools created by Live-SWE-agent. We compute the embedding for each tool based on the tool body (i.e., the content of the tool script) using a text embedding model (OpenAI text-embedding-3-small). To start off, we look at the types of tools created by Live-SWE-agent by categorizing them into common tool functionalities (e.g., edit, view, search, etc). We perform this categorization by using simple string matching based on the script filename of the tool. Figure 4(a) shows the visualization of the tools generated for SWE-bench Verified. We see that while there are definitely distinct clusters of common tools like edit, view, and search, there are still variations among them. For example, the edit tool cluster (highlighted in blue circles) is not a tightly concentrated dot, but instead is spreadout, showing that Live-SWE-agent can generate different editing tools depending on the problem. Furthermore, Live-SWE-agent also allows the agent to generate additional tools (the others category highlighted in red triangles). These are the more unique issue-specific tools such as a script that applies a very detailed patch to multiple files or a diff checking tool that shows the difference between two files. We also apply similar analysis to all studied datasets in Appendix B.
In addition to the types of tools created, we also examine how the tools created can change across different repositories and languages. Figure 4(b) shows the visualization of the tools created in SWE-Bench Pro categorized based on repositories. In SWE-Bench Pro, there are 11 different repositories, and not surprisingly, there are common tools created across different repositories (i.e., the overlapping clusters in the figure). On the other hand, there are also specific tools for certain repositories, for example we see a distinct cluster (highlighted in orange diamonds) for openlibrary, a codebase for cataloging literature. Since openlibrary deals with lots of raw data of a specialized format, the custom tools generated by Live-SWE-agent are specifically tailored for that. We also observe similar results when we break down the tools based on programming language of the repositories in Figure 4(c). Different problems, environments, and languages may require completely different specialized tools, highlighting again how using the same set of tools for all problems is suboptimal.
(d) Search tool
(e) Go file analyzer tool
Figure 5: Example custom tools
Effective and interesting custom tools. We now take a closer look at a few exemplary effective and interesting tools created by Live-SWE-agent. Figure 5 shows a custom search tool created by Live-SWE-agent. This example might appear simple and straightforward at first glance. However, it implements several important features: (1) supports focused searching within directories and pattern-based matching, (2) ignores irrelevant folders (__pycache__ and node_modules) and hidden folders (.git) during search, (3) displays relevant code context before and after the found code, and (4) limits results to top 20 matches to reduce context size. By baking in the exclusion of irrelevant folders in the search tool, agents do not have to add these folders as a separate flag during the searching process. Showing the relevant surrounding context is also helpful since oftentimes we want to find out how the code snippet of interest is being used. Furthermore, only showing the first 20 matches can be critically important to not drastically inflate the context window, allowing the agent to easily refine their search to limit results. An example bash command to emulate the functionality of the custom tool is shown at the top of Figure 5. We can see the high degree of complexity in terms of options, arguments, and flags used in this long command. On the other hand, we can simply call "python search_code.py ‘code’ src/" to perform the search via the custom tool. By only using basic bash commands, the agent will have to chain together multiple commands using complex flags with potential to incur errors (e.g., older grep versions do not support features like --exclude-dir) or substantial slowdowns. In fact, in our experimental results, we found that without access to custom tools, the agent may often take multiple steps to accomplish a basic task such as searching for relevant code. This issue is further amplified when steps like search may need to be done multiple times for one task thus drastically limiting the problem solving efficiency and effectiveness. By creating and introducing efficient and easy to use custom tools designed by agents themselves, Live-SWE-agent can improve agent performance on complex software engineering tasks.
Figure 5 shows another custom tool go_analyzer.py created by Live-SWE-agent. The tool can be used to analyze a Go file to find any struct and function definition, identifier references, and to obtain the imports used in the file. This example showcases the powerful generative capability of agents where the custom tool can be used as a simple static analyzer for the Go programming language to search and provide key information about a file. Different from the previous search tool example where one can technically mimic the behavior using bash commmands, abeit using very complex chain commands, the tool logic here is even more complex. The custom tool performs strategic pattern matching based on Go grammar and provides an interface to support multiple different usages. Furthermore, by saving this custom tool in an executable script, the agent can reuse the functionality multiple times. By creating and using this custom tool to better understand file structure and content, Live-SWE-agent is able to solve this issue333navidrome__navidrome-10108c63c9b5bdf2966ffb3239bbfd89683e37b7, SWE-Bench Pro that prior best performing baseline could not.
4.3 Ablation
To evaluate the effect of different setups of Live-SWE-agent, we use a random subset of 50 problems in SWE-bench Verified (see Appendix C.1 for the detailed list of problems). Unless otherwise stated, we use Claude 4.5 Sonnet in our ablation experiments.
Table 4: Ablation result across Live-SWE-agent setups on subset of SWE-bench Verified problems
Approach % Resolved # of tools created
Live-SWE-agent w/o tool creation 62.0% 0.00
Live-SWE-agent w/o reflection 64.0% 2.92
Live-SWE-agent 76.0% 3.28
Effectiveness of tool creation steps. Table 4 shows the impact of different components of Live-SWE-agent have on performance. We see that by removing the on-the-fly tool creation ability of Live-SWE-agent (i.e., using the base mini-SWE-agent), we achieve the lowest solve rate. The performance is improved when we indicate to the agent to create custom tools in the initial prompt (row Live-SWE-agent w/o reflection). However, the highest resolve rate is achieved once we explicitly ask the agent to reflect on the past trajectories to determine if they should create any tools after each step. In our experiments, we found that this reflection process provides a good reminder for the agent to create tools that are specifically designed for the particular issue. Furthermore, while the number of tools do not necessary correlate with solving performance, we also see that this reflection process creates more tools on average compared to only indicating the tool creation in the initial prompt. Live-SWE-agent is able to improve performance by reflecting on both the current problem and past trajectories to create custom tools on the fly.
Table 5: Ablation result across different LLM backends on subset of SWE-bench Verified problems
LLM mini-SWE-agent Live-SWE-agent
[Uncaptioned image] GPT-5-Nano 44.0% 14.0% ($\downarrow$-68.2%)
[Uncaptioned image] GPT-5-Mini 60.0% 58.0% ($\downarrow$-3.3%)
[Uncaptioned image] GPT-5 60.0% 68.0% ($\uparrow$13.3%)
Claude 3.7 Sonnet
46.0% 50.0% ($\uparrow$8.7%)
Claude 4 Sonnet
58.0% 64.0% ($\uparrow$10.3%)
Claude 4.5 Sonnet
62.0% 76.0% ($\uparrow$22.6%)
Different LLM backends. We now examine the effect of using different LLM backends for Live-SWE-agent. Table 5 compares the performance of the base mini-SWE-agent and Live-SWE-agent as we vary the LLM used on the same 50 problems used in the previous ablation experiment. We first observe that there are weaker LLMs for which using Live-SWE-agent to create custom tools on the fly even decreases performance. In particular, we see that for GPT-5-Nano, the results are significantly worse when using Live-SWE-agent compared to base mini-SWE-agent. After examining the trajectories, we found that GPT-5-Nano fails to understand the goal of creating custom tools and is often stuck in a loop thus leading to reduction in performance. This indicates that weaker LLMs may lack the high-level reasoning capabilities to synthesize useful tools on the fly. However, we see that Live-SWE-agent is able to improve more over the base mini-SWE-agent as we start to use more powerful models. Specifically, state-of-the-art LLMs such as Claude 4 Sonnet and GPT-5 achieves the highest relative resolve rate improvement when using Live-SWE-agent compared with mini-SWE-agent. This demonstrates the generalizability of Live-SWE-agent across high-performing LLMs and the potential for Live-SWE-agent to further improve performance especially as we continue to build more and more powerful LLMs.
Table 6: Result on a subset of SWE-bench Multilingual problems
Tool LLM % Resolved Avg. \$ Cost
mini-SWE-agent (yang2024sweagent)
Claude 4.5 Sonnet
40.0% \$0.59
Live-SWE-agent
Claude 4.5 Sonnet
46.0% \$0.66
SWE-bench Multilingual. In addition to evaluating on SWE-bench Verified and SWE-Bench Pro, we also test the generalizability of Live-SWE-agent on other benchmarks. We conducted an initial experiment on a subset of 50 problems (see Appendix C.2 for the detailed list of problems) in SWE-bench Multilingual (sbm), a benchmark of software engineering tasks across 9 programming languages (JavaScript, TypeScript, Rust, Ruby, Go, C/C++, PHP, and Java). Table˜6 shows the performance of Live-SWE-agent compared to mini-SWE-agent on SWE-bench Multilingual. We observed that Live-SWE-agent achieves a better performance by obtaining a resolve rate of 46.0%, while mini-SWE-agent only has a resolve rate of 40.0%. This demonstrates the generalizability of Live-SWE-agent on additional challenging problems.
4.4 Discussion and Future Work
Towards general, on the fly self-evolution. In this work, Live-SWE-agent primarily focuses on enabling agents to self-evolve through custom tool creation and usage. As discussed previously, the key idea of Live-SWE-agent, to improve and modify the agent itself on the fly, extends beyond creating new tools to also modifying the entire agent implementation. This includes the overall system prompt of the agent, how the agent interacts with the environment, and even the concrete workflow of the agent when it attempts to solve the issue. We believe that the entire agentic loop, much like any other software, can be modified and improved especially on the fly based on feedback and insights gained from solving a problem. Furthermore, we hope to extend the self-evolution loop across different tasks as well. Instead of discarding each evolved agent after a task is completed, we can save and serialize the useful tools and insights (via concepts like Skills skills) for future tasks. The agent can then load these useful tools and insights obtained while solving previous tasks on the fly to further improve performance and support continuous self-evolution across tasks.
Impact on LLM evaluation. Our experimental results on multiple benchmarks using multiple LLMs demonstrate that Live-SWE-agent can achieve state-of-the-art performance. This shows that Live-SWE-agent can provide an effective scaffold for unified evaluation of LLM performance on solving software development issues. Instead of using complex or proprietary agent scaffolds, Live-SWE-agent provides a simplistic and lightweight approach that can be easily added on top of any agent design to evaluate LLMs. Moreover, Live-SWE-agent not only evaluates the issue solving ability, but it can also test the tool-creation capability of LLMs and agents. Tools are one of the most important aspect of an agent and directly impacts the issue solving performance. The ability for LLMs to create custom tools, especially for solving complex software development issues, have not been extensively evaluated. To this end, we have already seen some interesting results in our evaluation where weaker LLMs lack high-level reasoning to create useful tools on the fly (see Section 4.3). Live-SWE-agent provides a unique framework to jointly evaluate both the tool creation and issue resolution ability of LLMs.
Applications beyond software issue resolution tasks. In addition to software issue resolution tasks, Live-SWE-agent can be easily applied to other challenging software engineering tasks like generating tests deng2023large, detecting and patching vulnerabilities lee2025sec, and synthesizing production-ready software from scratch zhao2024commit0. Compared with issue resolution tasks, other software problem domains will require even more task-specific and diverse tools and agent scaffolds. For example, to detect malicious vulnerabilities in commercial off-the-shelf binaries, we need binary analysis tools and decompilers. To optimize a large complex system, we need to apply profilers and tracing tools. Instead of using a fixed agent scaffold with basic tools that cannot adopt to different problem domains or painstakingly designing specialized agents for each individual task, Live-SWE-agent can easily generalize to solving tasks in different domains by automatically modifying itself on the fly based on the task at hand.
Self-evolution during LLM training. In this work, Live-SWE-agent is implemented as a lightweight modification to an existing agent, requiring no offline training or updates. However, the idea of on-the-fly self-evolution can also be easily extended to LLM training, where instead of learning from a fixed workflow and a static set of tools, the LLM learns to also create new tools and modify the scaffold itself during training. In this approach, the resulting LLM will have improved reasoning capabilities as the self-evolving training approach provides additional learning signals, allowing the LLM to solve more complex tasks. Through self-evolving training, LLMs can learn to better create useful, task-specific tools and modify advanced scaffolds dynamically based on the problem at hand. Moreover, the final trained LLM will be compatible with more runtime agent frameworks since it is able to learn from its own created tools and modified scaffolds rather than relying on predefined scaffolds. This adaptability allows it to be more robust and generalize effectively to new scaffold setups and even different tasks.
5 Related Work
5.1 Software Engineering Agents
Inspired by the human debugging process, where developers interact with the environmental feedback (like test failures) and learn from their earlier attempts, ChatRepair (xia2023conversational; xia2024automated) proposed the first interactive bug-fixing solution based on LLMs. Since ChatRepair, a large body of research work on bug fixing and general coding tasks aims to automatically provide LLMs more context information through multi-turn conversations (chen2023teaching; kong2025contrastrepair; yuan2024evaluating). More recently, foundation LLMs have seen substantial advances in their tool use and reasoning capabilities, and it becomes very natural to further equip feedback-driven solutions with such emergent LLM capabilities. In March 2024, Devin AI released the first AI software engineer, which can fully autonomously complete end-to-end software tasks, such as GitHub issue resolution (devin). The initial Devin release showed an impressive resolve rate of 13.86% on SWE-bench (jimenez2023swe), a dataset with thousands of real-world GitHub issues. Since then, a large number of dedicated software agent scaffolds have been proposed, including SWE-agent (yang2024swe), OpenHands (wang2024openhands), AutoCodeRover (zhang2024autocoderover), and Trae Agent (liu2024marscode). Such software agents typically equip the LLMs with a suite of coding tools and encourage the LLMs to autonomously decide the next actions for completing real-world software tasks. Different from the mainstream software agents, researchers have also proposed various AI software engineer solutions based on pre-defined workflows to challenge the necessity of complicated agent design, such as Agentless (xia2024agentless) and Moatless (moatless). Moreover, recent LLMs have increasingly been post-trained on massive real-world software data to better solve software engineering issues, including SWE-RL wei2025swerl, DeepSWE luodeepswe, DeepSeek V3.1 deepseekv31, MiniMax M1/M2 chen2025minimax, Kimi K2 team2025kimi, SWE-1.5 swe15, and Code World Model (CWM) carbonneaux2025cwm.
Due to the huge design space for software agents, it can be extremely challenging and costly to build an optimal agent scaffold. As a result, a number of self-improving software agents have been proposed very recently, including the Self-Improving Coding Agent (SICA) (robeyns2025sica), Darwin-Gödel Machine (DGM) (zhang2025darwin), and Huxley-Gödel Machine (HGM) (wang2025huxley). However, such self-improving agents require costly offline training on known benchmarks, and may not be generalize well across different LLMs, benchmarks, and issue types. Beyond the software engineering domain, prior work llmastool; wang2024voyager; qiu2025alitageneralistagentenabling; qian2024creatortoolcreationdisentangling; wang2024troveinducingverifiableefficient explored using LLMs to create tools for general reasoning or embodied tasks, but they do not target real-world software engineering problems. In contrast, in this paper, we propose Live-SWE-agent, the first live software agent capable of performing practical self-evolution on-the-fly during runtime when solving real-world issues. In this way, Live-SWE-agent requires no offline training at all and can be easily generalized to different LLMs and issue domains. Moreover, it also demonstrated superior performance compared to all existing self-improving software agents.
5.2 Benchmarks for Software Engineering Agents
To evaluate and demonstrate the performance of software engineering agents, a large number of datasets have been proposed. SWE-bench (jimenez2023swe) is one of the earliest and most widely used benchmark datasets for software agents. Besides the initial SWE-bench dataset, which includes thousands of real-world GitHub issues, researchers have also built curated subsets with high-quality and representative issues to support faster and more reliable evaluation, including SWE-bench Lite (jimenez2023swe) and SWE-bench Verified (sbv). Since SWE-bench mostly focused on Python projects, researchers have proposed a number of SWE-bench-style benchmarks for projects in multiple languages for more comprehensive agent evaluation, including SWE-PolyBench (rashid2025swe), SWE-bench Multilingual (sbm), and Multi-SWE-bench (zan2025multi). Also, SWE-bench relies heavily on manual effort for collecting benchmark instances and setting up executable environments, and is hardly scalable; as a result, researchers have also leveraged software agents to streamline issue collection and environment setup to build live and scalable benchmarks, including SWE-bench-Live (zhang2025swe) and SWE-rebench badertdinov2025swe. More recently, Scale AI has also constructed SWE-Bench Pro (deng2025swe), which aims to capture more realistic, complex, enterprise-level issues than SWE-bench. For rigorous evaluation of Live-SWE-agent, our study involved multiple widely used benchmarks, including SWE-bench Verified, SWE-bench Multilingual, SWE-bench Pro.
6 Conclusion
In this paper, we have proposed Live-SWE-agent, the first live software agent that can autonomously and continuously evolve itself on-the-fly during runtime when solving real-world software problems. More specifically, Live-SWE-agent starts with the most basic agent scaffold with only access to bash tools, and autonomously evolves its own scaffold implementation while solving real-world software problems. Our evaluation on the widely studied benchmarks (such as SWE-bench Verified and SWE-bench Pro) demonstrated that Live-SWE-agent outperforms state-of-the-art manually design software agents, demonstrating a promising future for live and self-evolving software agents.
Appendix A Additional experimental settings
We now describe additional experimental settings used by Live-SWE-agent for evaluation. As discussed in Section 3, we implement Live-SWE-agent on top of the mini-SWE-agent framework. We use the same default hyperparameter of mini-SWE-agent at the time of writing (maximum step limit of 250 and maximum cost of \$3 per issue).
For Gemini 3 Pro, we use temperature of 1 since it is recommended by the developers [^4] https://ai.google.dev/gemini-api/docs/gemini-3. For any Anthropic models in our experiments, we follow mini-SWE-agent and use a temperature of 0.0. For any OpenAI models in our experiments (e.g., GPT-5, GPT-5-Mini, GPT-5-Nano), since temperature 0.0 sampling is not supported, we use temperature of 1. To the best of our knowledge, this is also the case used in mini-SWE-agent when they use OpenAI models.
Appendix B Additional tool analysis
(a) Tools in SWE-bench Verified
(b) Tools in SWE-Bench Pro
Figure 6: Additional 2 dimensional t-SNE visualization of tools generated by Claude 4.5 Sonnet on SWE-bench Verified and SWE-Bench Pro. We label and display the embedding based on the repository name (Figure 6(a)) and tool name (Figure 6(b)). Note that for Figure 6(a), we only label two repositories in the legend due to space considerations. The two repositories were chosen as they have representative distinct clusters.
We perform additional tool analysis similar to the ones done in Section 4.2. We added the missing results including the visualization based on repository used for SWE-bench Verified and tool name for SWE-Bench Pro. Note that we did not include any results based on programming language for SWE-bench Verified since they are all repositories in Python. We observe similar patterns to the main evaluation.
To perform this analysis, we use sklearn implementation of PCA with n_components of 50. We then use the t-SNE implementation from sklearn with max_iter of 1000.
Appendix C Ablation problems
C.1 SWE-bench Verified ablation subset
Here we include the 50 randomly selected problems in SWE-bench Verified used in our ablation evaluations.
  • sympy__sympy-12489
  • django__django-16502
  • scikit-learn__scikit-learn-11310
  • django__django-13590
  • scikit-learn__scikit-learn-11578
  • matplotlib__matplotlib-21568
  • django__django-11276
  • django__django-11119
  • sympy__sympy-13031
  • sphinx-doc__sphinx-8459
  • django__django-14170
  • django__django-11066
  • sphinx-doc__sphinx-9658
  • django__django-14631
  • django__django-11749
  • scikit-learn__scikit-learn-13135
  • astropy__astropy-14369
  • matplotlib__matplotlib-24177
  • django__django-15252
  • pydata__xarray-7233
  • django__django-11815
  • sympy__sympy-18199
  • django__django-15467
  • psf__requests-5414
  • django__django-17084
  • scikit-learn__scikit-learn-15100
  • django__django-13925
  • django__django-11163
  • django__django-16595
  • django__django-16938
  • sympy__sympy-13877
  • django__django-16612
  • django__django-15629
  • pydata__xarray-3677
  • django__django-15503
  • psf__requests-1142
  • mwaskom__seaborn-3069
  • django__django-15127
  • astropy__astropy-13033
  • django__django-16429
  • django__django-16256
  • django__django-11133
  • django__django-13741
  • matplotlib__matplotlib-23476
  • scikit-learn__scikit-learn-12973
  • scikit-learn__scikit-learn-13779
  • sympy__sympy-23413
  • pytest-dev__pytest-7521
  • sympy__sympy-22456
  • django__django-16082
C.2 SWE-bench Multilingual ablation subset
We further evaluate on 50 randomly selected subset of problems in SWE-bench Multilingual. Below is the list of instance_ids of the problems in our subset.
  • apache__druid-13704
  • apache__druid-14136
  • apache__druid-15402
  • apache__lucene-13494
  • astral-sh__ruff-15543
  • axios__axios-6539
  • babel__babel-15445
  • burntsushi__ripgrep-2576
  • caddyserver__caddy-5404
  • caddyserver__caddy-5995
  • caddyserver__caddy-6051
  • facebook__docusaurus-9183
  • faker-ruby__faker-2970
  • fastlane__fastlane-20642
  • fluent__fluentd-3640
  • fmtlib__fmt-3750
  • javaparser__javaparser-4538
  • javaparser__javaparser-4561
  • jqlang__jq-2919
  • laravel__framework-48636
  • laravel__framework-52451
  • mrdoob__three.js-27395
  • nushell__nushell-12901
  • phpoffice__phpspreadsheet-3659
  • preactjs__preact-2757
  • preactjs__preact-2927
  • preactjs__preact-3562
  • preactjs__preact-3739
  • projectlombok__lombok-3326
  • reactivex__rxjava-7597
  • redis__redis-10068
  • redis__redis-10095
  • rubocop__rubocop-13503
  • rubocop__rubocop-13560
  • rubocop__rubocop-13627
  • rubocop__rubocop-13680
  • rubocop__rubocop-13705
  • sharkdp__bat-2650
  • sharkdp__bat-562
  • tokio-rs__axum-1730
  • tokio-rs__axum-2096
  • tokio-rs__axum-682
  • tokio-rs__tokio-6603
  • tokio-rs__tokio-6752
  • tokio-rs__tokio-6838
  • uutils__coreutils-6731
  • valkey-io__valkey-1842
  • valkey-io__valkey-790
  • vuejs__core-11739
  • vuejs__core-11870
Appendix D Prompts Used
Here we provide the detailed prompts used by Live-SWE-agent. Note that the majority of the prompt is constructed by mini-SWE-agent and we only make minimal changes to enable on-the-fly tool creation. Furthermore, we perform a slight modification: change the home directory indication of testbed to app in the initial prompt for all SWE-Bench Pro evaluations. This is done since the SWE-Bench Pro docker environments use app as the main source directory containing the repository code, different from SWE-bench environments. Apart from that, we use the same prompt for all our evaluations.
D.1 Initial prompt
Initial prompt for Live-SWE-agent <pr_description> Consider the following PR description: {{task}} </pr_description> <instructions> # Task Instructions ## Overview You’re a software engineer interacting continuously with a computer by submitting commands. You’$l$ be helping implement necessary changes to meet requirements in the PR description. {internallinenumbers*} Your task is specifically to make changes to non-test files in the current directory in order to fix the issue described in the PR description in a way that is general and consistent with the codebase. IMPORTANT: This is an interactive process where you will think and issue ONE command, see its result, then think and issue your next command. For each response: 1. Include a THOUGHT section explaining your reasoning and what you’re trying to accomplish 2. Provide exactly ONE bash command to execute ## Important Boundaries - MODIFY: Regular source code files in /testbed (this is the working directory for all your subsequent commands) - DO NOT MODIFY: Tests, configuration files (pyproject.toml, setup.cfg, etc.) ## Recommended Workflow 1. Analyze the codebase by finding and reading relevant files 2. Create a script to reproduce the issue 3. Edit the source code to resolve the issue 4. Verify your fix works by running your script again 5. Test edge cases to ensure your fix is robust ## Command Execution Rules You are operating in an environment where 1. You write a single command 2. The system executes that command in a subshell 3. You see the result 4. You write your next command Each response should include: 1. A **THOUGHT** section where you explain your reasoning and plan 2. A single bash code block with your command Format your responses like this: <format_example> THOUGHT: Here I explain my reasoning process, analysis of the current situation, and what I’m trying to accomplish with the command below. ```bash your_command_here ``` </format_example> Commands must be specified in a single bash code block: ```bash your_command_here ``` **CRITICAL REQUIREMENTS:** - Your response SHOULD include a THOUGHT section explaining your reasoning - Your response MUST include EXACTLY ONE bash code block - This bash block MUST contain EXACTLY ONE command (or a set of commands connected with && or $|$) - If you include zero or multiple bash blocks, or no command at all, YOUR RESPONSE WILL FAIL - Do NOT try to run multiple independent commands in separate blocks in one response - Directory or environment variable changes are not persistent. Every action is executed in a new subshell. - However, you can prefix any action with `MY_ENV_VAR=MY_VALUE cd /path/to/working/dir && ...` or write/load environment variables from files Example of a CORRECT response: <example_response> {internallinenumbers*} THOUGHT: I need to understand the structure of the repository first. Let me check what files are in the current directory to get a better understanding of the codebase. ```bash ls -la ``` </example_response> Example of an INCORRECT response: <example_response> THOUGHT: I need to examine the codebase and then look at a specific file. I’$l$ run multiple commands to do this. ```bash ls -la ``` Now I’$l$ read the file: ```bash cat file.txt ``` </example_response> If you need to run multiple commands, either: 1. Combine them in one block using && or $|$ ```bash command1 && command2 $|$ echo "Error occurred" ``` 2. Wait for the first command to complete, see its output, then issue the next command in your following response. ## Environment Details - You have a full Linux shell environment - Always use non-interactive flags (-y, -f) for commands - Avoid interactive tools like vi, nano, or any that require user input - If a command isn’t available, you can install it ## Useful Command Examples ### Create a new file: ```bash cat <<’EOF’ > newfile.py import numpy as np hello = "world" print(hello) EOF ``` ### View file content: ```bash # View specific lines with numbers nl -ba filename.py | sed -n ’10,20p’ ``` **IMPORTANT TOOL CREATION INSTRUCTIONS** ## Creating your own tools - You can also create your own tools in Python to help with your workflow - Compared to basic bash commands, the tools you create should be able to better aid your workflow in solving the task - Ensure each tool you create is in Python, contains informative outputs or error messages, and can be ran from the command line - You should at least create a simple edit tool that can help you effectively edit arbitrary files instead of using bash commands {internallinenumbers*} - The tools you create can be for any purpose, it does not need to be general, instead think about how it can help you specifically with the current task at hand ### Example of creating a custom tool: <example_response> THOUGHT: I noticed that in order to solve the issue I need to ... therefore I should create a custom tool to help me ... ```bash cat <<’EOF’ > /path/to/tool_name.py #!/usr/bin/env python3 import sys # Import other packages if needed def main(): # Your tool logic here ... if __name__ == "__main__": main() EOF ``` </example_response> ### Example of using the tool you created: <example_response> THOUGHT: Let me use the custom tool I created to help me with ... ```bash python /path/to/tool_name.py <<EOF your_input_here EOF ``` </example_response> ## Submission When you’ve completed your work (reading, editing, testing), and cannot make further progress issue exactly the following command: ```bash echo COMPLETE_TASK_AND_SUBMIT_FINAL_OUTPUT && git add -A && git diff --cached ``` This command will submit your work. You cannot continue working (reading, editing, testing) in any way on this task after submitting. </instructions>
Figure 7: The initial prompt for Live-SWE-agent, modified from the base mini-SWE-agent prompt.
D.2 Feedback message
Feedback message used by Live-SWE-agent <returncode>{{output.returncode}}</returncode> {% if output.output | length < 10000 -%} <output> {{ output.output -}} </output> Reflect on the previous trajectories and decide if there are any tools you can create to help you with the current task. Note that just because you can use basic bash commands doesn’t mean you should not create any tools that can still be helpful. {%- else -%} <warning> The output of your last command was too long. Please try a different command that produces less output. If you’re looking at a file you can try use head, tail, sed or create a tool to view a smaller number of lines selectively. If you’re using grep or find and it produced too much output, you can use a more selective search pattern. </warning> {%- set elided_chars = output.output | length - 10000 -%} <output_head> {{ output.output[:5000] }} </output_head> <elided_chars> {{ elided_chars }} characters elided </elided_chars> <output_tail> {{ output.output[-5000:] }} </output_tail> {%- endif -%}
Figure 8: The feedback message used by Live-SWE-agent after each agent step.